# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v3

@@ Abstract
Large language models (LLMs), trained on diverse data effectively acquire a breadth of information across various domains. However, their computational complexity, cost, and lack of transparency hinder their direct application for specialised tasks. In fields such as clinical research, acquiring expert annotations or prior knowledge about predictive models is often costly and time-consuming. This study proposes the use of LLMs to elicit expert prior distributions for predictive models. This approach also provides an alternative to in-context learning, where language models are tasked with making predictions directly. In this work, we compare LLM-elicited and uninformative priors, evaluate whether LLMs truthfully generate parameter distributions, and propose a model selection strategy for in-context learning and prior elicitation. Our findings show that LLM-elicited prior parameter distributions significantly reduce predictive error compared to uninformative priors in low-data settings. Applied to clinical problems, this translates to fewer required biological samples, lowering cost and resources. Prior elicitation also consistently outperforms and proves more reliable than in-context learning at a lower cost, making it a preferred alternative in our setting. We demonstrate the utility of this method across various use cases, including clinical applications. For infection prediction, using LLM-elicited priors reduced the number of required labels to achieve the same accuracy as an uninformative prior by 55%, 200 days earlier in the study.
@@ 1 Introduction
In fields such as clinical research, labelling data can be costly and time-consuming [1]. For example, in healthcare monitoring, patient data is recorded through wearables or remote monitoring devices, and observations are continuously collected. However, labelling such data for clinical outcomes, such as diagnosing illnesses or collecting and studying samples, can be challenging. In addition, estimating the uncertainty in clinical predictions is paramount [2], especially when available training data is sparse.
Expert knowledge of a problem domain can provide prior distributions of model parameters. This can improve a model’s performance when trained with few data points and enable estimates of uncertainties on parameters and predictions. However, expert knowledge about model parameter distributions is often difficult to obtain due to the challenges of performing prior elicitation from experts [3, 4], especially in low-resource settings.
Due to their diverse training data, language models offer an opportunity to begin resolving these issues. By considering a language model as an “expert” in a given field, we can extract insights about a predictive task that might lead to improved predictive models. This, and the abilities of language models to reason about problem settings outside of their intended use cases [5, 6, 7, 8, 9, 10], have inspired their use in performing prior elicitation [11, 12, 13]. Specifically, Selby et al. [12] show, by considering prior predictive distributions, that GPT-4 [10] can provide parameter distributions that allow for zero-shot performance (making predictions without training examples) of Bayesian models that compare to models trained on tens of data points. Furthermore, in Requeima et al. [13], the authors prompt language models with text and collected data points to elicit posterior predictive distributions for regression tasks, which they find outperforms Gaussian processes with radial basis function kernels on some tasks. This is done by repeatedly sampling a language model at a new feature value, given the previous input and output observations, and a given text description that contains prior information from a human expert. Given the rate of development of language models [14, 15, 16, 17], it is conceivable that their knowledge improves and broadens with time. This motivates the construction of a probabilistic framework to extract prior task knowledge from language models and evaluate the use of their elicited distributions for predictive tasks.
Figure 1: Prior elicitation using a language model.
In-context learning allows for the direct use of language models for predictive tasks. Here, language models are provided with task descriptions and demonstration examples from a training data set (in text form) and are asked to make predictions on unseen data [18, 19]. In this sense, in-context learning is similar to eliciting a prior distribution on a predictive model’s parameters and calculating its posterior predictive distribution, where the main differences arise from: (1) Language models are notoriously weak at performing arithmetic on numerical values [20]; (2) Language models lack transparency and may not be learning from demonstrations in a predictable or Bayesian way [21]; (3) Language models may not be able to faithfully provide their internal knowledge of a task when specifying a prior parameterisation for a separate predictive model or use that same prior knowledge in-context. (4) Language models incur a significant cost, and so their use for predictions in a high-frequency or low-resource setting is infeasible. To study the trade-offs between using in-context learning and prior elicitation, we aim to use the Bayes factor for model selection, and propose an approach to estimate a language model’s internal predictive prior and posterior models.
We are motivated by the following questions and make the following contributions:
  1. 1.
    Question: Can language models provide good prior distributions over the parameters of a predictive model, which enables the training of more robust models with less data?
    Method: We propose a method to elicit general prior parameter distributions for linear models from language models for predictive tasks (Section 2.2).
@@
    Finding: We find that using language models for expert prior elicitation provides better posterior accuracy and mean square error than an uninformative prior on a variety of numerical predictive tasks, enabling the construction of predictive models with fewer labelled examples (Section 3.1).
  2. 2.
    Question: How do the LLM’s in-context prior distributions compare to the ones we elicit?
    Method: We present an approach for extracting the in-context prior distribution (Section 2.3.2).
    Finding: We find that these distributions differ significantly from those elicited (Section 3.4).
  3. 3.
    Question: How can we check if the language model is performing Bayesian inference in-context?
    Method: By additionally providing the language model with training demonstrations, we extract the posterior distribution of the LLM’s in-context parameters (Section 2.3.2). We can then perform Monte Carlo sampling on the extracted prior and measure its difference to the extracted posterior.
    Finding: We find that the language model is inconsistent in its approximation of Bayesian inference and that its reliability depends on the task (Section 3.5).
  4. 4.
    Question: When is performance better with in-context learning?
    Method: We use the Bayes factor as a model selection method to decide between prior elicitation and in-context learning given a task (Section 2.4).
    Finding: Before considering cost and reliability, the Bayes factor depends on the task. However, after considering the in-context model’s inconsistent approximation of Bayesian inference and its cost, we conclude that prior elicitation is the preferred choice for all tasks tested (Section 3.6).
To implement our methods and reproduce our results, we provide details on the code [^1] https://github.com/alexcapstick/llm-elicited-priors in Appendix A.1 and complete experimental information in Appendix A.5.
@@ 2 Methods
2.1 Defining the task descriptions
The goal of this section is to define the space of prompts that we use to elicit priors and extract in-context predictions from a language model. We start by defining the space of all tasks as $\mathcal{T}$. An individual task, such as a specific classification or regression problem, will be referred to as $T\in\mathcal{T}$. Since a description of a task can be rephrased many times without changing its meaning, we must consider the space of possible task descriptions, referred to as $\mathcal{I}$. We say that the probability of a random text sequence $I\in\mathcal{I}$ being a relevant task description for task $T\in\mathcal{T}$ is: $p(I|T)$.
We assume that all relevant task descriptions have equal probability, that is, if $I_{i}$ and $I_{j}$ fully describe $T$, then $p(I_{i}|T)=p(I_{j}|T)$. We also assume that irrelevant task descriptions have $p(I|T)=0$. This means that for a task $T$, all of the mass of $p(\mathcal{I}|T)$ is uniformly spread across the relevant task descriptions only.
In reality, the number of relevant task descriptions would be too large to cover, since each word in the task description could be replaced with a synonym without losing the meaning of the full description. In Appendix A.11, we show that there are diminishing returns as you increase the number of ways of describing a task. For our experiments, we use $100$ relevant task descriptions to cover this space.
In practice, to produce $100$ task descriptions, we employ a language model to rephrase a single human written system role and user role $10$ times each and take their product. This means that using $100$ task descriptions is not particularly time-consuming and could be significantly increased by providing many base descriptions or prompting the language model to produce many more variants.
2.2 Using language models to elicit expert priors
2.2.1 Background
Bayesian inference allows us to incorporate our knowledge about a predictive task before observing the data. This is done using the prior knowledge $p(\theta)$ (over a predictive model parameterised by $\theta$), which can be our own intuition about which features are more important for a prediction and in what way they correlate. Additionally, using a prior allows us to report uncertainty over new predictions $(\tilde{x},\tilde{y})$, through the posterior predictive distribution $p(y|\tilde{x},D)$. If the prior is well-specified, using a prior can also reduce the number of data points required to construct a comparably accurate predictive model.
A key problem in machine learning is to define this prior distribution $p(\theta)$ given expert knowledge. Previous work studied how this distribution can be elicited from experts [3], which can be challenging and costly, or infeasible in low-resource settings. In parallel, research has shown that large language models (LLMs) encode knowledge about a surprising variety of domains [5, 6, 7, 8, 9, 10], which could be used to generate information about a predictive task [11, 12, 13], possibly supplemented with scientific discussions between experts and the language model[22].
Choi et al. [11] explores the use of an LLM to perform feature selection, causal discovery, and to modify the reward function in a reinforcement learning task. In the first case, they observe that when listing features from two datasets, GPT-3 can correctly identify the features that correspond to a single specified predictive task. Additionally, when applied to causal discovery, the language model is often able to predict the directionality of causal relationships in the Tüebingen Cause-Effect Pairs dataset [23]. This inspires us to explore the possibility of using LLMs for eliciting prior parameter distributions over linear models.
Further, Selby et al. [12] elicit parameters of a beta distribution to define a prior for predictive tasks where results from human experts are available. The priors from four LLMs were visually compared to those elicited from human experts using the SHELF [3] method, which demonstrated significant differences between experts and LLMs. In our work, we extend this by using the LLM to elicit multiple Gaussian distributions as priors. We combine these to create a mixture model for a more general distribution over our parameters and test their utility for linear predictive models.
Furthermore, Zhu and Griffiths [24] describe the use of iterated learning to elicit prior distributions from language models for causal learning, proportion estimation, and predicting everyday quantities. The authors employ a Markov chain Monte Carlo method to access an LLM’s prior knowledge by performing successive inference using previous model outputs that tend towards a limiting distribution. In doing this, the authors assume that in-context learning is a form of Bayesian inference. However, in their work Zhu and Griffiths [24] do not consider how prior elicitation from LLMs can be used for predictive models or test whether an LLM is performing Bayesian inference, both of which we study in this work.
2.2.2 Our contribution: LLM-elicited priors
Our first contribution, therefore, focuses on proposing a method for eliciting an LLM’s prior knowledge of a task to construct general distributions on the parameters of a predictive model that improve performance with less data compared to a predictive model with an uninformative prior and produce uncertainty about its predictions. An overview of this process is given in Figure 1.
Formally, given a language model $M$ and task $T$, we will obtain a single Gaussian prior for each feature, for each task description $I$. This is done by sampling $(\mu_{k},\sigma_{k})$ from $p_{M,T}(\mu,\sigma|I_{k})$, where we obtain a single $\mu_{k}$ and $\sigma_{k}$ by prompting the language model to provide a guess of the mean and standard deviation of a Gaussian prior for each feature on a linear model.
@@ 2.2.2 Our contribution: LLM-elicited priors
By taking a mixture of these individual Gaussian distributions, we construct a prior $p(\theta|M,T)=p_{M,T}(\theta)$ over our model parameters $\theta$:
@@
In our experiments, our mixture has over $100$ components, using several different task descriptions. Mixtures of Gaussians are able to approximate any probability density with enough components, allowing us to approximate the language model’s belief about the task, even if non-Gaussian. We would also expect the language model to have learnt how to parameterise a Gaussian distribution from its training, making them good candidates. However, mixtures of other distributions can be used instead and may be more useful for particular tasks.
@@
$p_{M,T}(y|\tilde{x},\mathcal{D})=\int_{\Theta}\sum_{k=1}^{K}p_{M,T}(y|\tilde{x},\theta)p_{M,T}(\theta|\mathcal{D},I_{k})p_{M,T}(I_{k})d\theta$
Here, $p_{M,T}(I)$ refers to the probability distribution discussed in Section 2.1. The distribution $p_{M,T}(\theta|\mathcal{D},I)$ describes the posterior distribution over the parameterisation of a predictive model, which is calculated using the available data $\mathcal{D}$ and the prior parameterisation we extract from the language model.
The practical sampling of these distributions is discussed in Section 3.1 and Appendix A.5.2.
Since we are using a language model’s prior knowledge and a given training dataset to make predictions, this method is an alternative to in-context learning. The difference is that prior elicitation uses a marginal distribution and Bayesian inference to address some of the shortfalls of in-context learning.
@@ 2.3 In-context learning
2.3.1 Background
In-context learning aims to use the knowledge that language models acquire during training to make predictions on unseen data. Specifically, this method requires providing a language model with a description of the predictive task and some demonstration examples that allow it to learn “in-context” before the user queries the predictions on some unseen data.
Probabilistically, given a language model $M$ and task $T$, and many task descriptions $I_{k}\in\mathcal{I}$, a prediction of $\tilde{y}$ on features $\tilde{x}$ using in-context learning with observed data $\mathcal{D}$ is:
We extend this by introducing the concept of an internal model that the LLM is updating and then using to make its predictions in-context, parameterised by $\varphi$ for each task description:
$p_{M,T}(y|\tilde{x},\mathcal{D})\\ =\int_{\Phi}\sum_{I\in\mathcal{I}}p_{M,T}(y|\tilde{x},\mathcal{D},\varphi)p_{M,T}(\varphi|\mathcal{D},I)p_{M,T}(I)d\varphi$
Where $\varphi$ is itself described by a prior distribution $p_{M,T}(\varphi|I)$, a posterior distribution $p_{M,T}(\varphi|\mathcal{D},I)$, and is unseen. This $\varphi$ is unseen because when we perform in-context learning on a language model for a task description $I$, we only receive $\tilde{y}\sim p_{M,T}(y|\tilde{x},\mathcal{D},I)$ in return.
The inference used to sample $\tilde{y}$ can be performed multiple times to build a sampling distribution that approximates the posterior predictive distribution on $\tilde{y}$ given $\tilde{x}$ and $\mathcal{D}$: $p_{M,T}(\tilde{y}|\tilde{x},\mathcal{D})$.
In the above notation, it is assumed that the output of the language model contains only the predicted label $\tilde{y}$. Alternatively, the log-likelihood of the tokens representing the possible labels can be measured from the language model’s output, leading to a modified notation as in Dong et al. [19].
Learning from demonstrations in-context has proven to be effective for several natural language tasks, demonstrating the surprising ability of language models to perform outside of their intended use [18, 19, 25, 26, 15]. In Brown et al. [18], the authors study GPT-3’s strong ability on natural language tasks after seeing a few examples, one example, or none at all. This work inspired others to study the extent to which language models can learn from demonstrations given as text alongside a task description. Further, Xie et al. [27] study the mechanics of in-context learning and describes it from a Bayesian Inference perspective. They find that in-context learning succeeds when language models can infer a shared concept across individual examples in a prompt. Min et al. [21] argue that in-context learning allows the language model to observe the label space, the distribution of input text, and the task format, as they find that swapping the real labels in demonstrations for random labels does not significantly reduce accuracy. Most of the existing work focuses on natural language tasks [19], leaving numerical problems less explored and a focus of our work.
In Hollmann et al. [28], the authors train a transformer model to solve synthetically generated classification tasks from a prior over tabular datasets. Given a new task and demonstrations, this transformer performs in-context learning to make predictions using inference steps only. Furthermore, Akyürek et al. [29] and Von Oswald et al. [30] discuss the types of learning algorithms that language models (and more generally, transformers) can implement in-context. Focusing on linear models, Akyürek et al. [29] show that, by construction, transformers are capable of performing gradient descent and closed-form ridge regression. Akyürek et al. [29] also demonstrate that trained in-context learners closely match predictors trained through gradient descent. Further, Von Oswald et al. [30] show that transformers, by incorporating multilayer perceptrons into their architectures, can also solve non-linear regression tasks, and Dalal and Misra [31] propose a Bayesian framework to reason about a language model’s internal state and suggest that in-context learning is consistent with Bayesian learning. In our work, we test empirically, through the approximation of the internal model used by a language model for its in-context predictions, whether for a given task, the language model is performing Bayesian inference before and after seeing demonstrations.
@@ 2.3.1 Background
To fully compare in-context learning and prior elicitation, we must consider two main questions: (1) Does the language model use the same prior in-context that we elicit? (2) Does in-context learning approximate Bayesian inference? To answer these questions, we aim to extract, in the case of a linear model, the in-context predictive model’s prior and posterior parameter distributions.
2.3.2 Our contribution: Extracting the in-context prior and posterior
If we assume that the class of the internal predictive model used by the language model can be specified through a prompt (Appendix A.3), then we can estimate the distribution on the parameters of the internal in-context model, $p_{M,T}(\varphi|\mathcal{D})$ and $p_{M,T}(\varphi)$, through the sampling distribution of a maximum likelihood estimator.
@@ 2.3.2 Our contribution: Extracting the in-context prior and posterior
By specifying the structure of the internal predictive model to be used by the language model $M$ within our task descriptions $I$ for a task $T$, we can write the internal model as: $f_{\varphi}:X\rightarrow Y$ where $f$ represents the class of the predictive model, $\varphi$ is the random variable associated with its parameters, and $X$ and $Y$ correspond to the domain and range of the predictive model, respectively. Depending on the model class $f$ corresponds to, $y\in Y$ could be a probability distribution over a set of possible targets or the prediction of the target itself.
@@
Performing inference on our language model (when it is using its internal predictive model for in-context predictions) returns a $y$ given some $x$. Therefore, although we do not have access to $\varphi$ directly, we do have access to $p_{M,T}(y|x,\varphi)$ and we can perform maximum likelihood estimation to generate a sampling distribution for $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$.
@@
Given a language model $M$ and task $T$, we construct the sample distribution (following Murphy [32]) as follows:
@@
Where $\text{MLE}_{M,T}(f_{\varphi}(X)|X,I)$ refers to the maximum likelihood estimate of the parameters $\varphi$ given input feature $X$ and in-context predictions $f_{\varphi}(X)$ for a task description $I$. This is a sample of the estimated prior distribution over the parameters $\varphi$.
We can also provide $\mathcal{D}$ examples with the task description to get posterior samples:

### v4

# AutoElicit: Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

@@ Abstract
Large language models (LLMs) acquire a breadth of information across various domains. However, their computational complexity, cost, and lack of transparency often hinder their direct application for predictive tasks where privacy and interpretability are paramount. In fields such as healthcare, biology, and finance, specialised and interpretable linear models still hold considerable value. In such domains, labelled data may be scarce or expensive to obtain. Well-specified prior distributions over model parameters can reduce the sample complexity of learning through Bayesian inference; however, eliciting expert priors can be time-consuming. We therefore introduce AutoElicit to extract knowledge from LLMs and construct priors for predictive models. We show these priors are informative and can be refined using natural language. We perform a careful study contrasting AutoElicit with in-context learning and demonstrate how to perform model selection between the two methods. We find that AutoElicit yields priors that can substantially reduce error over uninformative priors, using fewer labels, and consistently outperform in-context learning. We show that AutoElicit saves over 6 months of labelling effort when building a new predictive model for urinary tract infections from sensor recordings of people living with dementia.
@@ 1 Introduction
Labelling data in medicine can be a costly and time-consuming undertaking (Ghassemi et al., 2020; Cooper et al., 2023). We are particularly interested in the development of automated tools to assist with healthcare monitoring for people living with dementia. In our study, participant data is recorded through passive monitoring devices. Although input features are continuously collected at scale, the process of labelling adverse health conditions, such as urinary tract infection (UTI), can be challenging. Expert knowledge of a task, as privileged information (Vapnik et al., 2015), can provide prior distributions of model parameters, which often reduces the number of samples needed to learn performant models while also enabling more accurate estimates of predictive uncertainty. However, this expert knowledge about model parameters can be difficult to obtain due to the challenges of performing expert prior elicitation, particularly in low-resource settings (Mikkola et al., 2021).
Linear models are widely used in real-world applications, especially in healthcare and as part of our wider study on dementia care, due to their interpretability, robustness, and clinical familiarity (Tennant et al., 2024; Capstick et al., 2024; Ramírez Medina et al., 2025). Many studies do not use prior distributions on their models, possibly due to the challenges associated with elicitation, but would likely benefit greatly from them.
Due to their diverse training data and rate of development  (Naveed et al., 2023), large language models (LLMs) offer a new opportunity to fully or partially automate the prior elicitation process, especially given recent advances in reasoning methods (Huang & Chang, 2023). By prompting an LLM to approximate a specific form of expertise in a given field, it is possible to extract insights about a predictive model’s prior based on the LLM’s training data. This motivates the construction of a probabilistic framework to extract prior task knowledge from language models in the form of distributions over the parameters of a linear predictive model. AutoElicit combines developments in LLMs with trusted and robust linear predictive models, that are already widely understood and utilised in real-world applications.
Figure 1: AutoElicit: Prior elicitation using a language model. The left hand figure demonstrates the process, whilst the right hand figure illustrates the benefits of using AutoElicit, achieving the same peak accuracy after $n\approx 15$ labels, 220 days earlier in the study.
In-context learning has been studied as a way to use LLMs for predictive tasks (Dong et al., 2022). Here, LLMs are given a task description and demonstration examples and asked to make predictions on unseen data directly (Brown et al., 2020). There are several reasons why this prompted prediction may not suffice. First, LLMs can incur a large computational cost to run, rendering their use in high-frequency or low-resource settings unlikely. Second, applications may require arithmetic operations within models, an area where LLMs continue to struggle (Spithourakis & Riedel, 2018). Third, LLMs, unless explicitly trained on a specific predictive problem, lack transparency and may not perform Bayesian inference (Min et al., 2022), which can lead to unreliable posterior updates and predictions.
Our work makes the following contributions:
@@
  • We introduce AutoElicit, an algorithm to generate LLM-elicited priors over linear models. We find using LLMs for expert prior elicitation (Figure 1) provide higher posterior accuracy and lower mean square error than an uninformative prior on various numerical predictive tasks, improving models with fewer examples.
  • To compare prior elicitation with in-context learning (ICL), we develop an approach to extract the implicit priors used in ICL which we find to be less informative. To further understand the differences, we study whether an LLM performs Bayesian inference during ICL. We extract the in-context posterior distribution and measure its divergence from posterior Monte Carlo samples of the extracted prior. We find that the language model inconsistently approximates Bayesian inference on real-world datasets.
  • We study when ICL is better than AutoElicit and show how to use the Bayes factor to guide model selection, finding that LLM-elicited priors are the preferred choice for all tasks tested. When predicting UTIs in our study on dementia care, we find that AutoElicit priors result in performant accuracy over six months earlier than with uninformative priors (Figure 1).
Our results provide insight on when and how to use LLMs to build better linear models, with fewer samples. Open-source code [^1] https://github.com/alexcapstick/llm-elicited-priors to reproduce our work is described in Appendix A.1.
2 Related Work
Garthwaite et al. (2005) and O’Hagan et al. (2006) discuss eliciting prior distributions from human experts using a standardised procedure: Experts are trained in probabilities to provide judgements of a task, which can be used to fit probability distributions. Later, Oakley & O’Hagan (2019) provide software and an e-learning course, which trains human experts in the knowledge required to describe informative priors. However, this can be challenging, costly, or infeasible in low-resource settings (Mikkola et al., 2021).
In parallel, research shows that LLMs encode varied domain knowledge, which could be extracted for predictive tasks, possibly in discussions with experts (Gao et al., 2024). Recent works have used LLMs for prior elicitation in multiple ways. In data pre-processing, Choi et al. (2022) and Hollmann et al. (2024) show that LLMs can perform feature selection and engineering respectively, whilst Zhang et al. (2023) and Zhang et al. (2024) use LLMs to specify a modelling pipeline given a task description. This suggests language models might also be able to provide information over the parameters of predictive models directly. In this direction, Gouk & Gao (2024) use LLMs to generate synthetic data from public datasets, which define priors on linear models through a separate likelihood term. However, they do not consider that these datasets might have appeared in the LLM’s training (Appendix A.8), making it challenging to evaluate. Related, Selby et al. (2024) evaluate LLMs for estimating correlations between features and targets, and Zhu & Griffiths (2024) use them to elicit priors for causal learning, proportion estimation, and everyday quantities. However, this knowledge is not used for predictive models. Further related work is given in Appendix A.2.1. In contrast to the literature, we provide a framework for using LLMs to elicit priors over the parameters of linear models directly, which we evaluate on public and private predictive tasks.
LLMs can also be used to produce posterior predictive distributions. In-context learning is an active research area, mainly focused on natural language tasks (Dong et al., 2022), with many works aiming to explain the underlying mechanisms. Relevant to us, Requeima et al. (2024) prompt LLMs with task descriptions and data to generate posterior predictive distributions for regression tasks. Additionally, Hollmann et al. (2022) train a transformer to solve synthetically generated tasks from a prior on tabular datasets and use in-context learning at prediction time. However, these methods sample LLMs for every prediction, which is resource-intensive and potentially unreliable. Additional related work is given in Appendix A.2.2. In our work, we propose a strategy to approximate the in-context prior and posterior distributions when imitating a linear model, which we compare to LLM-elicited priors. Using these, we empirically evaluate whether in-context learning approximates Bayesian inference before using model selection.
LLMs have been evaluated for use in healthcare (Thirunavukarasu et al., 2023), demonstrating a broad knowledge of clinical problems and strong performance on medical questions, even without fine-tuning (Nori et al., 2023; Singhal et al., 2023). Despite this, their direct use for decision making, through in-context learning, is not recommended as they sometimes produce inaccurate information. However, linear models have found wide and trusted uses in clinical applications (Ramírez Medina et al., 2025; Tennant et al., 2024), since their model weights have intuitive descriptions. This motivates us to combine LLMs, which can reason about clinical problems and risk factors, with linear models, which are interpretable and familiar to clinicians.
3 Background
3.1 Prior elicitation with experts
Bayesian inference incorporates knowledge about a predictive task before data is observed (Murphy, 2022). A prior $p(\theta)$ is placed on a predictive model, which holds intuitions about the parameters $\theta$. Well-specified priors allow us to estimate predictive uncertainty through the posterior predictive distribution and reduce the number of data points needed to construct comparably accurate predictive models.
Prior elicitation constructs prior distributions on model parameters, typically through expert judgement (Oakley & O’Hagan, 2019; Gosling, 2018). This can be resource intensive and requires access to one or more domain experts who have been trained in prior elicitation (Mikkola et al., 2021).
3.2 Linear models
Because of their simplicity, linear models have interpretable parameter weights, making them trusted in clinical contexts. Model weights are explicitly understood in their effect on predictions, as they multiply directly with features, making them ideal to use with prior elicitation in healthcare.
3.3 In-context learning
In-context learning uses the knowledge that LLMs acquire during training to make predictions on unseen data. Here, LLMs attempt to learn a function for predicting unseen data based on a task description and demonstration examples.
Let the space of all predictive tasks be $\mathcal{T}$, a single classification or regression task be $T\in\mathcal{T}$, and the space of possible LLM prompts be $\mathcal{I}$. Given an LLM $M$ and $K$ task descriptions $I_{k}\sim p(I|T)=p_{T}(I)$, a prediction of $\tilde{y}$ on features $\tilde{x}$ using in-context learning with training data $\mathcal{D}$ is:
Although the mechanisms of in-context learning are under active research, we denote its implicit parameters as $\varphi$, implied by a task description and dataset:
$p_{M,T}(y|\tilde{x},\mathcal{D})\\ =\int_{\Phi}\sum_{k=1}^{K}p_{M,T}(y|\tilde{x},\varphi)p_{M,T}(\varphi|\mathcal{D},I_{k})p_{M,T}(I_{k})d\varphi$ (1)
Here $\varphi$ is itself described by a prior distribution $p_{M,T}(\varphi|I_{k})$ and a posterior distribution $p_{M,T}(\varphi|\mathcal{D},I_{k})$. $\varphi$ is latent, since in-context learning only returns $\tilde{y}\sim p_{M,T}(y|\tilde{x},\mathcal{D},I)$.
@@ 4 Methods
4.1 AutoElicit: Using language models to elicit priors
We aim to make expert prior elicitation more accessible by using the knowledge of LLMs in robust and clinically trusted predictive models. This motivates us to develop a framework for eliciting general prior distributions for linear models from LLMs, AutoElicit (Figure 1), which often have both technical knowledge of probability distributions and domain knowledge of a task. Our goal is to improve model performance with less data compared to an uninformative prior, whilst allowing model parameters to be interpretable.
Given an LLM $M$ and task $T$, we obtain a single Gaussian prior for each feature, for each task description $I_{k}$ by asking the LLM to provide its best guess of the mean and standard deviation: $(\mu_{k},\sigma_{k})\sim p_{M,T}(\mu,\sigma|I_{k})$. Here, we use $K=100$ task descriptions, produced by asking an LLM to rephrase one human-written “system” and “user” role $10$ times and taking their product (Appendix A.7.3).
@@ 4.1 AutoElicit: Using language models to elicit priors
Taking a mixture over the task descriptions, we construct a prior $p_{M,T}(\theta)$ over linear model parameters $\theta$:
@@
Mixtures of Gaussians approximate any distribution with enough components, allowing us to fully capture the LLM’s prior belief. We also expect that LLMs have learnt how to parametrise Gaussian distributions, although our framework does allow for other mixtures. In our work, this mixture has $100$ components, using as many task descriptions, and in Appendix A.12 we find diminishing differences in the LLM-elicited priors as we increase the number of descriptions.
@@
$p_{M,T}(y|\tilde{x},\mathcal{D})\\ =\int_{\Theta}\sum_{k=1}^{K}p_{M,T}(y|\tilde{x},\theta)p_{M,T}(\theta|\mathcal{D},I_{k})p_{M,T}(I_{k})d\theta$ (3)
Where $p_{M,T}(I)$ describes the distribution over task descriptions and $p_{M,T}(\theta|\mathcal{D},I)$ describes the posterior distribution over the parameterisation of a predictive model, which is calculated using the available data $\mathcal{D}$ and the AutoElicit prior over $\theta$.
This $p_{M,T}(\theta)$ is the prior that we use for our linear models and is an alternative to one elicited from experts. However, additionally including expert information $E$ in the prompts can provide yet more informative priors: $p_{M,T}(\theta|E)$. This way, AutoElicit also supplements a typical prior elicitation process but only requires experts to provide knowledge as natural language, likely needing less training and cost. This approach enables the use of an LLM’s knowledge in new environments, with more controls and transparency.
The natural comparison for AutoElicit priors is an uninformative prior. However, since we are using an LLM’s prior knowledge and training data to make new predictions, this is also an alternative to in-context learning. The difference being that prior elicitation uses a marginal distribution and Bayesian inference, which address some of the shortfalls of in-context learning on numerical predictive tasks.
@@ 4.2 In-context learning
To compare in-context learning and AutoElicit, we must consider: (1) Does the LLM use the same elicited prior in-context? and (2) Does in-context learning approximate Bayesian inference? To answer these, we extract the prior and posterior distributions of the in-context model.
Extracting the in-context prior and posterior
@@
We can write the implicit in-context model as: $f_{\varphi}:X\rightarrow Y$ where $f$ represents the class of the predictive model, which we specify in the prompt, $\varphi$ is the random variable associated with its parameters, and $X$ and $Y$ correspond to its domain and range. Based on the model class $f$ corresponds to, $y\in Y$ will either be a predicted target or its probability.
@@
In-context learning returns a $y$ given an $x$. Since we do not have access to $\varphi$ directly, we propose to use $p_{M,T}(y|x,\varphi)$ and maximum likelihood estimation (MLE) to generate a sampling distribution for $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$.
@@
To do this, given a language model $M$ and task $T$, we construct the sample distribution (Murphy, 2022):
@@
Where $\text{MLE}_{M,T}(f_{\varphi}(X)|X,I)$ refers to the MLE of the parameters $\varphi$ given input $X$, in-context predictions $f_{\varphi}(X)$, and task description $I$. This provides a sample of the estimated prior on $\varphi$ for the domain of $X$. Further, additionally providing training data $\mathcal{D}$, returns posterior samples: